# Understanding uncertainty in Bayesian cluster analysis

### v1

@@ Appendix D Experiments: additional material
For scenario (1), we investigated the impact of cluster separation on posterior uncertainty using datasets where the model is correctly specified but the mixture components have varying degrees of overlap. Each dataset consists of $n=600$ observations in two dimensions, drawn from a mixture of four equally weighted Gaussian components. The component means are symmetrically positioned at $(m,m)$, $(m,-m)$, $(-m,-m)$, and $(-m,m)$, where the separation parameter $m$ (with identity covariance matrix) controls the distance between clusters. The value of $m$ was varied across a dense grid to explore a wide range of overlap, specifically $m\in\{0.1,~{}0.3,~{}0.5,~{}0.7,~{}0.9,~{}1.1,~{}1.3,~{}1.5,~{}1.7,~{}1.9,~{}2.0,~{}2.2,~{}2.4,~{}2.6,$ 2.8,3.0}2.8,~{}3.0\}2.8 , 3.0 }. For each value of $m$, ten independent datasets were generated with different random seeds, with points assigned to each component with equal probability. A graphical representation is provided in Figure 11.
@@
[…]
@@
Figure 11: Illustration for the data generating process for scenario (1) in the synthetic experiments. The data is generated from a four-components mixture Gaussian distribution, with each component centered around the points $(x_{i},y_{i})$ for $i=1,\ldots,4$, with $|x_{i}|=|y_{i}|=m$. The covariance matrix for each component is diagonal with marginal variances equal to 1. Several datasets are displayed for different values of the center location parameter $m$.
D.1 Posterior Compression

### v2

@@ D.2 Examples in Section 4
For scenario (1), we investigated the impact of cluster separation on posterior uncertainty using datasets where the model is correctly specified but the mixture components have varying degrees of overlap. Each dataset consists of $n=600$ observations in two dimensions, drawn from a mixture of four equally weighted Gaussian components. The component means are symmetrically positioned at $(m,m)$, (m,m)(m,-m), (m,m)(-m,-m), and (m,m)(-m,m), where the separation parameter $m$ (with identity covariance matrix) controls the distance between clusters. The value of $m$ was varied across a dense grid to explore a wide range of overlap, specifically $m\in\{0.1,~0.3,~0.5,~0.7,~0.9,~1.1,~1.3,~1.5,~1.7,~1.9,~2.0,~2.2,~2.4,~2.6,$ 2.8,3.0}2.8,~3.0\}. For each value of $m$, ten independent datasets were generated with different random seeds, with points assigned to each component with equal probability. A graphical representation is provided in Figure 14.
@@
Specifically, for setting (a) we consider Gaussian distributions truncated to the unit square $0,1]^{2}$. For each standard deviation $\sigma$ in the set $\{0.1,\,0.2,\,0.4,\,0.6,\,0.8,\,1,\,3\}$, data are generated from a normal distribution centered at $(0.5,0.5)$ with covariance matrix $\sigma^{2}I$. Points falling outside the unit square are discarded and resampled until exactly $n$ points are retained. For small $\sigma$, the data is concentrated near the center; as $\sigma$ increases, the resulting distribution approaches a uniform distribution over the unit square.
@@
Figure 14: Illustration for the data generating process for scenario (1) in the synthetic experiments. The data is generated from a four-components mixture Gaussian distribution, with each component centered around the points $(x_{i},y_{i})$ for $i=1,\ldots,4$, with $|x_{i}|=|y_{i}|=m$. The covariance matrix for each component is diagonal with marginal variances equal to 1. Three datasets are displayed for different values of the center location parameter $m$.
D.3 Posterior compression
@@
Appendix E Real data analysis
E.1 Investigating projection patterns at the single-neuron resolution
Figure 16: Heatmap of the empirical projection strength (computed as the neuron’s barcode counts divided by its total count), with rows representing neurons and columns corresponding to the 11 target areas.
The auditory cortex data (Chen et al., 2019) illustrated in Figure 16 suggests diversity in the projections of neurons. HBMAP (Wade et al., 2025) employs a hierarchical Bayesian mixture of Dirichlet-multinomials to directly model the count data, effectively integrate data across brains, and identify projection patterns. In the MCMC algorithm of HBMAP, we combine five parallel runs with the number of iterations to 20,000, burn-in of 10,000 and thinning of 5, resulting in 10,000 posterior samples of partitions. The marginal posterior on the number of clusters is concentrated around 22-32; the optimal clustering estimate contains 27, 37, and 50 clusters under the VI, ARI, and Binder’s loss; and the posterior similarity matrix suggests both stability but also uncertainty in clusters (see Figure 17).
[…]
(a) minVI partition
(b) Posterior similarity matrix
Figure 17: Neurons belonging to different clusters of the minVI partition are separated by dashed lines in the heatmap of the data in Panel 17(a) and by solid lines in the posterior similarity matrix in Panel 17(b).
(a) WASABI elbow plot
(b) Weight and number of clusters in the particles
Figure 18: The elbow plot of the Wasserstein distance against the number of particles suggests $L=3$ particles (Panel 18(a)). The weight and number of clusters in each of the particles is shown in Panel 18(b).
[…]
(a) Stable meet clusters with low group contribution to the EVI
[…]